A leather jacket costs a clothing store manager
Php 14 000. Find the selling price of the leather
jacket if the mark up rate is 40%.

Answer by ikleyn(52786)

.
A leather jacket costs a clothing store manager
Php 14 000. Find the selling price of the leather
jacket if the mark up rate is 40%.
~~~~~~~~~~~~~~~

Then the selling price is 14000 + 0.4*14000 = (1+0.4)*14000 = 1.4*14000 = Php 19600.    ANSWER

Solved.